# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v3

@@ Abstract
ABACUS (Atomic-orbital Based Ab-initio Computation at USTC) is an open-source software for first-principles electronic structure calculations and molecular dynamics simulations. It mainly features density functional theory (DFT) and molecular dynamics functions and is compatible with both plane-wave basis sets and numerical atomic orbital basis sets. ABACUS serves as a platform that facilitates the integration of various electronic structure methods, such as Kohn-Sham DFT, stochastic DFT, orbital-free DFT, and real-time time-dependent DFT, etc. In addition, with the aid of high-performance computing, ABACUS is designed to perform efficiently and provide massive amounts of first-principles data for generating general-purpose machine learning potentials, such as DPA models. Furthermore, ABACUS serves as an electronic structure platform that interfaces with several AI-assisted algorithms and packages, such as DeePKS-kit, DeePMD, DP-GEN, DeepH, DeePTB, etc.
@@ I Introduction
The theoretical foundations of density-based methodologies can be traced to the Thomas-Fermi (TF) model proposed in 1927. The TF model established the first kinetic energy functional based on the uniform electron gas approximation Thomas (1927); Fermi (1927). Although this orbital-free (OF) framework pioneered the conceptual basis for density functional theory (DFT), its oversimplified treatment of kinetic energy limited chemical accuracy Wang and Carter (2002). In 1964, Hohenberg and Kohn Hohenberg and Kohn (1964) established two theorems. First, the external potential of a many-electron system is uniquely determined by its electron density. Second, a universal energy functional exists whose minimum value corresponds to the ground-state electron density. Later in 1965, Kohn and Sham Kohn and Sham (1965) developed the Kohn-Sham density functional theory (KS-DFT), whose accuracy is determined by the exchange-correlation functional. This framework allows a non-interacting electron system where the kinetic energy can be computed exactly via electronic wavefunctions. Crucially, the non-interacting system can yield the electron density of the real system via the so-called self-consistent field (SCF) method.
@@
Over the past few decades, driven by developments of the DFT method itself and continuous refinement of numerical algorithms, various popular KS-DFT packages emerge Blaha et al. (2020); Zhang et al. (2020a); Romero et al. (2020); Nakata et al. (2020); Kühne et al. (2020); Tancogne-Dejean et al. (2020); Prentice et al. (2020); Sun et al. (2020); Giannozzi et al. (2020); Garc´ıa et al. (2020); Kresse and Joubert (1999); Frisch et al. (2016). In addition, with the exponential growth in computing capabilities, the KS-DFT method has become a cornerstone to predict various properties of materials in physics, chemistry, materials science, biology, geophysics, etc. Despite these achievements, challenges still remain. For example, the community increasingly demands DFT implementations that combine enhanced computational efficiency with improved hardware adaptability across different architectures. Recently, with rapid advancements in artificial intelligence (AI) technologies, an important question arises: can these new technologies drive substantial improvements in computational accuracy and efficiency for electronic structure methods such as KS-DFT and OF-DFT?
@@
In this review paper, we introduce the ABACUS package, which was initiated around 2006 and has achieved a series of progress over the past two decades Chen, Guo, and He (2010a, 2011a); Li et al. (2016); Lin, Ren, and He (2020); Zheng, Ren, and He (2021); Lin, Ren, and He (2021a, b); Qu et al. (2022a, b); Liu and Chen (2022a); Dai, Jin, and He (2022); Liu et al. (2023); Lin et al. (2024). For instance, the ABACUS package supports both plane-wave (PW) basis and numerical atomic orbitals (NAO) basis for first-principles electronic structure calculations, geometry relaxation and molecular dynamics simulations. In addition, ABACUS offers robust support for highly efficient parallelization of diverse electronic structure methods, leveraging MPI, OpenMP, and CUDA to optimize performance across distributed and accelerated computing architectures. Based on recent advancements in AI technologies for atomistic simulations and electronic structure methods Zhang et al. (2018a); Wang et al. (2018); Chen et al. (2021, 2022a); Kirkpatrick et al. (2021); Jia et al. (2020); Han, Zhang, and Weinan (2019); Chen et al. (2020); Wen et al. (2022), researchers have identified significant potential for AI to transform both electronic structure methods and their corresponding software implementations. In this context, the effort to create an integrated platform that integrates physics-based models, high-performance computing, and AI capabilities has become particularly critical. Since 2021, the ABACUS development team has maintained an ongoing collaboration with the DeepModeling open-source community aba (2024).
@@
As illustrated in Fig. 1, ABACUS aims to establish an electronic structure algorithmic platform and computational engine tailored for the AI for Science (AI4S) paradigm. This integration features three aspects. First, state-of-the-art electronic structure method developments with AI-assisted algorithms. Second, molecular dynamics simulations to provide data for the OpenLAM project ope (2024a), which is designed to generate a general-purpose machine learning model for the periodic table. Third, the platform can be potentially used in various industry fields such as battery, catalyst, electronics, ferroelectricity, magnetic materials, high-entropy alloys, etc.
@@
In this work, we provide a detailed overview of the latest developments in ABACUS. Sec. II details implementation of general features in ABACUS, such as the pseudopotentials, exchange-correlation functionals, SCF method, geometry relaxation and molecular dynamics, etc. Sec. III describes the implementation of the KS-DFT, stochastic DFT (sDFT), and OF-DFT methods with the plane wave basis. Moreover, recent advances for AI-assisted kinetic energy density functionals (KEDFs) are introduced for OFDFT. Sec. IV introduces the linear combination of atomic orbitals (LCAO) calculation methods, where numerical atomic orbitals are served as basis to solve the KS equation and yield forces and stress, as well as newly developed methods such as AI-assisted electronic structure method DeePKS, DFT+U, real-time time-dependent DFT (RT-TDDFT) and hybrid functionals, all of which are based on the numerical atomic orbital basis. Sec. V mainly focuses on AI-assisted molecular dynamics methods and describes the roled played by ABACUS. Sec. VI presents various software interfaces between ABACUS and other packages. Sec. V.1 introduces the pseudopotential and numerical atomic orbital library. Finally, we conclude in sec. VII.
@@ II.1 Kohn-Sham Equation
The KS equation can be solved via different types of basis sets, such as plane waves Kresse and Joubert (1999); Giannozzi et al. (2020); Romero et al. (2020), numerical atomic orbitals Garc´ıa et al. (2020); Boker et al. (2011); Nakata et al. (2020); Blum et al. (2009), real-space grids Chelikowsky, Troullier, and Saad (1994); Tancogne-Dejean et al. (2020), Gaussian orbitals Frisch et al. (2016); Kühne et al. (2020), wavelet basis sets Ratcliff et al. (2020), etc. Currently, ABACUS supports both plane waves and NAOs as basis sets. On the one hand, the plane wave basis can accurately and efficiently handle condensed matter systems with a relatively small number of atoms. On the other hand, the NAOs are centered on atoms and the required number of basis is substantially less than the plane wave basis. Therefore, the NAO basis is suitable to simulate systems containing thousands of atoms Kühne et al. (2020); Garc´ıa et al. (2020).
@@ II.1.1 Pseudopotentials
The Coulomb potential exhibits a singularity at the atomic nucleus, necessitating an extensive basis set for the precise representation of “near-core” wavefunctions, which imposes substantial computational overhead. To overcome this challenge, several pseudopotential methodologies have been developed in the past few decades Kleinman (1980); Bachelet and Schlüter (1982); Rappe et al. (1990); Vanderbilt (1990); Troullier and Martins (1991); Kresse and Hafner (1994); Kresse and Joubert (1999); Hamann (2013). These pseudopotentials achieve significant reduction in computational costs while preserving computational accuracy.
ABACUS supports both norm-conserving pseudopotentials (NCPPs) Hamann, Schlüter, and Chiang (1979) and ultrasoft pseudopotentials (USPPs) Vanderbilt (1990). Specifically, the pseudopotential operator $\hat{V}_{\rm ps}$ is decomposed into local and nonlocal components in NCPPs as
@@
where the local pseudopotential $\hat{V}_{\rm local}$ equals the all-electron functions beyond a cutoff radius. Furthermore, $\hat{V}_{\mathrm{KB}}$ is composed of a set of separable nonlocal projectors as proposed by Kleinman and Bylander Kleinman and Bylander (1982), lated extended by Hamann Hamann (2013). The formula is
@@
Here $\beta_{lm}$ is a projector with angular quantum number $l$ and magnetic quantum number $m$, the coefficients are in the $D_{ll^{\prime}}$ matrix. In addition, these projects $\beta_{lm}$ are expressed as the product of a radial function and a spherical harmonic function. In cases such as transition metals or magnetic systems, the non-linear core corrections Louie, Froyen, and Cohen (1982) are usually adopted in the pseudopotential file to improve the transferability of pseudopotentials. For the spin-orbital coupling effect, it can be incorporated into the non-local projectors in the $\hat{V}_{\mathrm{KB}}$ term Theurich and Hill (2001); Corso and Conte (2005); Cuadrado et al. (2021).
Compared to NCPPs, the Ultrasoft pseudopotentials Vanderbilt (1990) offer substantially improved computational efficiency by removing the norm-conservation constraint, which allows greater flexibility in constructing the pseudo-wavefunction within the core region, yielding inherently softer wavefunctions that reduce computational overhead. As a result, USPPs facilitate faster convergence of basis sets expansions, though this advanrage comes with the trade-off of requiring more sophisticated implementation with KS-DFT packages.
@@ II.1.2 Exchange-Correlation Functionals
Since the invention of KS-DFT, numerous density functional approximations have been developed. The earliest implementation, the local density approximation (LDA) Kohn and Sham (1965), is expressed as a functional of the spin-summed electron density $\rho(\mathbf{r})$. Subsequent refinements introduced the local spin density approximation (LSDA) Von Barth and Hedin (1972); Vosko, Wilk, and Nusair (1980), which incorporates spin density components $\rho^{\uparrow}$ and $\rho^{\downarrow}$ within a broken-symmetry Slater determinant framework, thereby correctly describing the dissociation limit of H2.
@@
Further advancement was achieved with meta-GGA functionals, which incorporate the kinetic energy density $\tau(\mathbf{r})$. This class of functionals, exemplified by the strongly constrained and appropriately normed (SCAN) functional Sun, Ruzsinszky, and Perdew (2015), offers enhanced accuracy over GGAs. Variants of SCAN, such as rSCAN Bartók and Yates (2019) and R2SCAN Furness et al. (2020), were subsequently developed to improve numerical stability and convergence while largely retaining its accuracy.
One may also include a fraction of exact Hartree-Fock exchange energy into the exchange-correlation term, and the resulting functional is called hybrid functionals, such as the HSE06 Heyd, Scuseria, and Ernzerhof (2003); Ge and Ernzerhof (2006), PBE0 Adamo and Barone (1999a) and B3LYP functionals Stephens et al. (1994).
Analogous to spin-unpolarized systems, approximations for $E_{\rm xc}[\rho^{\uparrow},\rho^{\downarrow}]$ are typically constructed hierarchically. The simplest approach, the LSDA Miehlich et al. (1989), expresses $E_{\rm xc}$ solely as a function of the local spin densities $\rho^{\sigma}(\mathbf{r})$. More refined treatments, such as the generalized gradient approximation (GGA), introduce additional dependence on the gradients of the spin densities $\nabla\rho^{\sigma}(\mathbf{r})$. Further sophistication is achieved in meta-GGA functionals through the incorporation of the kinetic energy density, while hybrid functionals explicitly include fractional exact exchange integrals.
The LDA, GGA, and meta-GGA functionals have already been implemented in ABACUS for both plane-wave and numerical atomic orbital basis sets. Both LDA and GGA functionals are accessible through either native implementations or the Libxc library Lib (2024), whereas the meta-GGA functional is currently restricted to Libxc. Benchmark studies of SCAN, rSCAN Bartók and Yates (2019), and r2SCAN Furness et al. (2020) within ABACUS are detailed in Ref. Liu et al., 2023. Efficient hybrid functional calculations are available for localized basis sets, utilizing combined capabilities of Libxc and the local resolution of identity (LRI) methodology (see Sec. IV.4 for computational details).
@@ II.1.3 Non-Collinear Spin
Here $\mathbf{b}_{xc}$ is the effective exchange-correlation magnetic field, defined as the functional derivative of the exchange-correlation energy with respect to the magnetization density. During Hamiltonian construction, the potential must be reconstructed in its spinor form to operate on the two-component Kohn-Sham wavefunctions.
@@ II.2 Self-Consistent Field Method
Figure 3: ABACUS performs self-consistent field (SCF) calculations with two available basis sets: PW and NAOs. The PW basis set employs iterative diagonalization methods like CG and Davidson to solve the Kohn-Sham equations. For NAOs, it uses exact diagonalization or the low-scaling method PEXSI Lin et al. (2009, 2013). ABACUS applies different mixing algorithms for various types of calculations, including non-magnetic, collinear, and non-collinear. Additionally, ABACUS supports the density matrix mixing method for DFT+U, hybrid functionals, and DeePKS calculations with the NAO basis set.

### v4

@@ Abstract
ABACUS (Atomic-orbital Based Ab-initio Computation at USTC) is an open-source software for first-principles electronic structure calculations and molecular dynamics simulations. It mainly features density functional theory (DFT) and molecular dynamics functions and is compatible with both plane-wave basis sets and numerical atomic orbital basis sets. ABACUS serves as a platform that facilitates the integration of various electronic structure methods, such as Kohn-Sham DFT, stochastic DFT, orbital-free DFT, and real-time time-dependent DFT, etc. In addition, with the aid of high-performance computing, ABACUS is designed to perform efficiently and provide massive amounts of first-principles data for generating general-purpose machine learning potentials, such as DPA models. Furthermore, ABACUS serves as an electronic structure platform that interfaces with several AI-assisted algorithms and packages, such as DeePKS-kit, DeePMD, DP-GEN, DeepH, DeePTB, HamGNN, etc.
@@ I Introduction
The theoretical foundations of density-based methodologies can be traced to the Thomas-Fermi (TF) model proposed in 1927. The TF model established the first kinetic energy functional based on the uniform electron gas approximation. Thomas (1927); Fermi (1927) Although this orbital-free (OF) framework pioneered the conceptual basis for density functional theory (DFT), its oversimplified treatment of kinetic energy limited chemical accuracy. Wang and Carter (2002) In 1964, Hohenberg and Kohn Hohenberg and Kohn (1964) established two theorems. First, the external potential of a many-electron system is uniquely determined by its electron density. Second, a universal energy functional exists whose minimum value corresponds to the ground-state electron density. Later in 1965, Kohn and Sham Kohn and Sham (1965) developed the Kohn-Sham density functional theory (KS-DFT), whose accuracy is determined by the exchange-correlation functional. This framework allows a non-interacting electron system where the kinetic energy can be computed exactly via electronic wavefunctions. Crucially, the non-interacting system can yield the electron density of the real system via the so-called self-consistent field (SCF) method.
@@
Over the past few decades, driven by developments of the DFT method itself and continuous refinement of numerical algorithms, various popular KS-DFT packages emerge. Blaha et al. (2020); Zhang et al. (2020a); Romero et al. (2020); Nakata et al. (2020); Kühne et al. (2020); Tancogne-Dejean et al. (2020); Prentice et al. (2020); Sun et al. (2020); Giannozzi et al. (2020); García et al. (2020); Kresse and Joubert (1999); Frisch et al. (2016) In addition, with the exponential growth in computing capabilities, the KS-DFT method has become a cornerstone to predict various properties of materials in physics, chemistry, materials science, biology, geophysics, etc. Despite these achievements, challenges still remain. For example, the community increasingly demands DFT implementations that combine enhanced computational efficiency with improved hardware adaptability across different architectures. Recently, with rapid advancements in artificial intelligence (AI) technologies, an important question arises: can these new technologies drive substantial improvements in computational accuracy and efficiency for electronic structure methods such as KS-DFT and OF-DFT?
@@
In this review paper, we introduce the ABACUS package, which was initiated around 2006 and has achieved a series of progress over the past two decades. Chen, Guo, and He (2010a, 2011a); Li et al. (2016); Lin, Ren, and He (2020); Zheng, Ren, and He (2021); Lin, Ren, and He (2021a, b); Qu et al. (2022a, b); Liu and Chen (2022a); Dai, Jin, and He (2022); Liu et al. (2023); Lin et al. (2024) For instance, the ABACUS package supports both plane-wave (PW) basis and numerical atomic orbitals (NAO) basis for first-principles electronic structure calculations, geometry relaxation and molecular dynamics simulations. In addition, ABACUS offers robust support for highly efficient parallelization of diverse electronic structure methods, leveraging MPI, OpenMP, and CUDA to optimize performance across distributed and accelerated computing architectures. Based on recent advancements in AI technologies for atomistic simulations and electronic structure methods, Zhang et al. (2018a); Wang et al. (2018); Chen et al. (2021, 2022a); Kirkpatrick et al. (2021); Jia et al. (2020); Han, Zhang, and Weinan (2019); Chen et al. (2020); Wen et al. (2022) researchers have identified significant potential for AI to transform both electronic structure methods and their corresponding software implementations. In this context, the effort to create an integrated platform that integrates physics-based models, high-performance computing, and AI capabilities has become particularly critical. Since 2021, the ABACUS development team has maintained an ongoing collaboration with the DeepModeling open-source community. aba (2024)
@@
As illustrated in Fig. 1, ABACUS aims to establish an electronic structure algorithmic platform and computational engine tailored for the AI for Science (AI4S) paradigm. This integration features three aspects. First, state-of-the-art electronic structure method developments with AI-assisted algorithms. Second, molecular dynamics simulations to provide data for the OpenLAM project, ope (2024a) which is designed to generate a general-purpose machine learning model for the periodic table. Third, the platform can be potentially used in various industry fields such as battery, catalyst, electronics, ferroelectricity, magnetic materials, high-entropy alloys, etc.
@@
In this work, we provide a detailed overview of the latest developments in ABACUS. Sec. II details implementation of general features in ABACUS, such as the pseudopotentials, exchange-correlation functionals, SCF method, geometry relaxation and molecular dynamics, etc. Sec. III describes the implementation of the KS-DFT, stochastic DFT (sDFT), and OF-DFT methods with the plane wave basis. Moreover, recent advances for AI-assisted kinetic energy density functionals (KEDFs) are introduced for OFDFT. Sec. IV introduces the linear combination of atomic orbitals (LCAO) calculation methods, where numerical atomic orbitals are served as basis to solve the KS equation and yield forces and stress, as well as newly developed methods such as hybrid functional, AI-assisted electronic structure method DeePKS, DFT+U and real-time time-dependent DFT (RT-TDDFT), all of which are based on the numerical atomic orbital basis. Sec. V mainly focuses on machine-learning-based interatomic potentials and describes the role played by ABACUS. Sec. VI presents various software interfaces between ABACUS and other packages. Finally, we conclude in sec. VII.
@@ II.1 Kohn-Sham Equation
The KS equation can be solved via different types of basis sets, such as plane waves, Kresse and Joubert (1999); Giannozzi et al. (2020); Romero et al. (2020) numerical atomic orbitals, García et al. (2020); Boker et al. (2011); Nakata et al. (2020); Blum et al. (2009) real-space grids, Chelikowsky, Troullier, and Saad (1994); Tancogne-Dejean et al. (2020) Gaussian orbitals, Frisch et al. (2016); Kühne et al. (2020) wavelet basis sets, Ratcliff et al. (2020) etc. Currently, ABACUS supports both plane waves and NAOs as basis sets. On the one hand, the plane wave basis can accurately and efficiently handle periodic systems with a relatively small number of atoms. On the other hand, the NAOs are centered on atoms and the required number of basis is substantially less than the plane wave basis. Therefore, the NAO basis is suitable to simulate systems containing thousands of atoms. Kühne et al. (2020); García et al. (2020)
@@ II.1.1 Pseudopotentials
The Coulomb potential exhibits a singularity at the atomic nucleus, necessitating an extensive basis set for the precise representation of “near-core” wavefunctions, which imposes substantial computational overhead. To overcome this challenge, several pseudopotential methodologies have been developed in the past few decades. Kleinman (1980); Bachelet and Schlüter (1982); Rappe et al. (1990); Vanderbilt (1990); Troullier and Martins (1991); Kresse and Hafner (1994); Kresse and Joubert (1999); Hamann (2013) These pseudopotentials achieve significant reduction in computational costs while preserving computational accuracy.
ABACUS supports both norm-conserving pseudopotentials (NCPPs) Hamann, Schlüter, and Chiang (1979) and ultrasoft pseudopotentials (USPPs). Vanderbilt (1990) Specifically, the pseudopotential operator $\hat{V}_{\rm ps}$ is decomposed into local and nonlocal components in NCPPs as
@@
where the local pseudopotential $\hat{V}_{\rm local}$ equals the all-electron functions beyond a cutoff radius. Furthermore, $\hat{V}_{\mathrm{KB}}$ is composed of a set of separable nonlocal projectors as proposed by Kleinman and Bylander, Kleinman and Bylander (1982) lated extended by Hamann. Hamann (2013) The formula is
@@
Here $\beta_{lm}$ is a projector with angular quantum number $l$ and magnetic quantum number $m$, and its weighting factor is $D_{ll^{\prime}}$. In addition, these projectors $\beta_{lm}$ are expressed as the product of a radial function and a spherical harmonic function. In cases such as transition metals or magnetic systems, the non-linear core corrections Louie, Froyen, and Cohen (1982) are usually adopted in the pseudopotential file to improve the transferability of pseudopotentials. For the spin-orbital coupling effect, it can be incorporated into the non-local projectors in the $\hat{V}_{\mathrm{KB}}$ term. Theurich and Hill (2001); Corso and Conte (2005); Cuadrado et al. (2021)
Compared to NCPPs, the Ultrasoft pseudopotentials Vanderbilt (1990) offer substantially improved computational efficiency by removing the norm-conservation constraint, which allows greater flexibility in constructing the pseudo-wavefunction within the core region, yielding inherently softer wavefunctions that reduce computational overhead. As a result, USPPs facilitate faster convergence of basis set expansions, though this advantage comes with the trade-off of requiring more sophisticated implementation within KS-DFT packages.
@@ II.1.2 Exchange-Correlation Functionals
Since the invention of KS-DFT, numerous density functional approximations have been developed. The earliest implementation, the local density approximation (LDA), Kohn and Sham (1965) is expressed as a functional of the spin-summed electron density $\rho(\mathbf{r})$. Subsequent refinements introduced the local spin density approximation (LSDA), Von Barth and Hedin (1972); Vosko, Wilk, and Nusair (1980) which incorporates spin density components $\rho^{\uparrow}$ and $\rho^{\downarrow}$ within a broken-symmetry Slater determinant framework, thereby correctly describing the dissociation limit of H2.
@@
Further advancement was achieved with meta-GGA functionals, which incorporate the kinetic energy density $\tau(\mathbf{r})$. This class of functionals, exemplified by the strongly constrained and appropriately normed (SCAN) functional, Sun, Ruzsinszky, and Perdew (2015) offers enhanced accuracy over GGAs. Variants of SCAN, such as rSCAN Bartók and Yates (2019) and R2SCAN, Furness et al. (2020) were subsequently developed to improve numerical stability and convergence while largely retaining its accuracy.
One may also include a fraction of exact Hartree-Fock exchange energy into the exchange-correlation term, and the resulting functional is called hybrid functionals, such as the HSE06, Heyd, Scuseria, and Ernzerhof (2003); Ge and Ernzerhof (2006) PBE0 Adamo and Barone (1999a) and B3LYP functionals. Stephens et al. (1994)
Analogous to spin-unpolarized systems, approximations for $E_{\rm xc}[\rho^{\uparrow},\rho^{\downarrow}]$ are typically constructed hierarchically. The simplest approach, the LSDA, Miehlich et al. (1989) expresses $E_{\rm xc}$ solely as a function of the local spin densities $\rho^{\sigma}(\mathbf{r})$. More refined treatments, such as the generalized gradient approximation (GGA), introduce additional dependence on the gradients of the spin densities $\nabla\rho^{\sigma}(\mathbf{r})$. Further sophistication is achieved in meta-GGA functionals through the incorporation of the kinetic energy density, while hybrid functionals explicitly include fractional exact exchange integrals.
The LDA, GGA, and meta-GGA functionals have already been implemented in ABACUS for both plane-wave and numerical atomic orbital basis sets. Both LDA and GGA functionals are accessible through either native implementations or the Libxc library, Lib (2024) whereas the meta-GGA functional is currently restricted to Libxc. Benchmark studies of SCAN, rSCAN, Bartók and Yates (2019) and r2SCAN Furness et al. (2020) within ABACUS are detailed in Ref. Liu et al., 2023. Efficient hybrid functional calculations are available for localized basis sets, utilizing combined capabilities of Libxc and the local resolution of identity (LRI) methodology (see Sec. IV.4 for computational details).
@@ II.1.3 Non-Collinear Spin
Here $\mathbf{b}_{xc}$ is the effective exchange-correlation magnetic field, defined as the functional derivative of the exchange-correlation energy with respect to the magnetization density, and it can be decomposed into three independent components ($\mathbf{b}^{x}_{xc},\mathbf{b}^{y}_{xc},\mathbf{b}^{z}_{xc}$). During Hamiltonian construction, the potential must be reconstructed in its spinor form to operate on the two-component Kohn-Sham wavefunctions.
@@ II.2 Self-Consistent Field Method
Figure 3: ABACUS performs self-consistent field (SCF) calculations with two available basis sets: PW and NAOs. The PW basis set employs iterative diagonalization methods like CG and Davidson to solve the Kohn-Sham equations. For NAOs, it uses exact diagonalization or the low-scaling method PEXSI. Lin et al. (2009, 2013) ABACUS applies different mixing algorithms for various types of calculations, including non-magnetic, collinear, and non-collinear. Additionally, ABACUS supports the density matrix mixing method for DFT+U, hybrid functionals (EXX), and DeePKS calculations with the NAO basis set.